OpenStudy (naga447):

A bag contains 1 blue, 2 white, and 2 red marbles. match the probability with the term. A) 10% B) 16% C) 8% D) 20% 1) the probability of red and red, with replacement 2)the probability of white and white, without replacement 3) the probability of blue and red, with replacement 4) the probability of white and red, without replacement

OpenStudy (anonymous):

1) red and red with replacement = (2/5)*(2/5) = 0.16 = 16% 2) white and white without replacement = (2/5)*(1/4) = 0.1 = 10% 3) blue and red with replacement = (1/5)*(2/5) = 0.08 = 8% 4) white and red without replacement = (2/5)*(2/4) = 0.2 = 20%

OpenStudy (naga447):

Thanks!!!